# Accelerating Multimodal Large Language Models via Dynamic Visual-Token Exit and the Empirical Findings

### v1

@@ Abstract
The excessive use of visual tokens in existing Multimoal Large Language Models (MLLMs) often exhibits obvious redundancy and brings in prohibitively expensive computation. To gain insights into this problem, we first conduct extensive empirical studies on the attention behaviors of MLLMs, and summarize three main inference stages in MLLMs: (i) Early fusion between tokens is first accomplished quickly. (ii) Intra-modality modeling then comes to play. (iii) Multimodal reasoning resumes and lasts until the end of inference. In particular, we reveal that visual tokens will stop contributing to reasoning when the text tokens receive enough image information, yielding obvious visual redundancy. Based on these generalized observations, we propose a simple yet effective method to improve the efficiency of MLLMs, termed dynamic visual-token exit (DyVTE). DyVTE uses lightweight hyper-networks to perceive the text token status and decide the removal of all visual tokens after a certain layer, thereby addressing the observed visual redundancy. To validate VTE, we apply it to a set of MLLMs, including LLaVA, VILA, Eagle and InternVL, and conduct extensive experiments on a bunch of benchmarks. The experiment results not only show the effectiveness of our VTE in improving MLLMs’ efficiency, but also yield the general modeling patterns of MLLMs, well facilitating the in-depth understanding of MLLMs. Our code is anonymously released at https://github.com/DoubtedSteam/DyVTE.
@@ 1 Introduction
Figure 1: Illustration of three main stages observed in MLLMs. During inference, an MLLM will gradually go through three main stages, i.e., early fusion between all tokens, intra modality modeling of the same-modal tokens, and multimodal reasoning between all tokens again. When the text tokens receive enough image information, visual tokens will stop contributing to inference regardless of their activities. Based on this observation, we propose a novel dynmiac visual-token exit (DyVTE) method for MLLMs.
@@
Recently, the rapid development of vision-language learning has been witnessed with the great success of large language models (LLMs) [50, 19, 2, 57, 49, 6, 1, 28]. Numerous efforts are devoted to equipping LLMs with general multimodal capability, i.e., multimodal large language models (MLLMs) [40, 33, 3, 53, 12]. To overcome visual shortcoming [46, 60], recent MLLMs often resort to higher-resolution images as input, which is often accompanied with a multitude of visual tokens [33, 45, 53, 12].
@@
Despite effective, the excessive use of visual tokens also leads to prohibitively expensive computation. For instance, compared with LLaVA 1.5 [40], LLaVA-NeXT [33] adopts about 1728 more visual tokens, resulting in about 4 times FLOPs. Moreover, recent progresses also show that these visual tokens often exist heavy redundancy in MLLMs. For instance, Ye et al. [59] show that randomly dropping half visual tokens barely affects the performance on common VL tasks, such as MMB [41] and SQA [42]. In this case, the efficient use of visual tokens has recently become a research hot-spot, attracting an influx of attention [9, 54, 16, 52, 7].
@@
To gain insights into visual redundancy in MLLMs, we first conduct extensive empirical studies about the attention behaviors of MLLMs, and summarize three main stages of their inference, as illustrated in Fig.1. At the first stage, an MLLM quickly accomplishes the exchange of multimodal information at its shallow layers, which we term (i) early fusion. Afterwards, this information exchange fades away, and the self-attention between tokens of the same modality becomes the main activity in the second stage, termed (ii) intra-modality modeling. At the last stage, the visual tokens will resume the information propagation to the text tokens, and this process will continue until the end of inference, and we term it (iii) multimodal reasoning. Through extensive comparisons, we reveal that these observed behaviors of MLLMs are generalized across different models [40, 37, 45, 11] and tasks [26, 27, 46, 60, 41], as shown in our experimental section.
Behind these general behaviors of MLLMs, we also observe interesting findings directly related to visual redundancy. As summarized above, the visual tokens will continue to diffuse their semantics to the text ones at the multimodal reasoning stage, while its self-modeling also often plays at the same time. However, we find that this cross-modal interaction in fact barely contributes to multimodal reasoning after some layers. In other words, multimodal reasoning basically acts within text tokens when the text tokens receive enough visual semantics. For instance, dropping all visual tokens at some top layers of MLLMs barely impedes the prediction and also makes not obvious changes to the self-attention patterns of text tokens, as shown in our experimental section. Overall, these results imply that visual tokens can in fact exit at a certain layer of MLLMs without affecting the following reasoning.
@@
Motivated by these observations, we propose a simple yet effective method for improving MLLMs’ efficiency in this paper, termed dynamic visual-token exit (DyVTE) as depicted in Fig.2. In particular, we apply lightweight hyper-networks to perceive text token status and decide the removal of all visual tokens, thereby improving efficiency and speeding up inference. Compared with previous efforts like visual token pruning [4, 8], we study the efficiency issue of MLLMs from a new perspective, i.e., perceiving the learning status of text tokens about multimodal reasoning rather than evaluating the redundancy of specific visual tokens. Moreover, these two strategies are orthogonal and can work collaboratively, as proven in our experiments. In addition, DyVTE also greatly differs from common dynamic early exiting methods [10, 21, 48, 25], which often refers to a inference break, i.e., skipping the rest layers for a direct prediction. In contrast, text tokens will continue to transform in DyVTE, which better meets our empirical findings.
@@
To validate DyVTE, we apply it to a set of advanced MLLMs with varying scales, including LLaVA-1.5 [40], VILA [37], Eagle [45] and InternVL [11], on a bunch of widely-used VL and MLLM benchmarks [26, 27, 46, 35, 24, 60]. The experimental results show that our DyVTE can greatly reduce the computation overhead of MLLMs, while still maintaining their competitive performance on various benchmarks. For instance, our DyVTE reduces the computation overhead of LLaVA-1.5 by up to 45.7% without significant performance drops. Moreover, DyVTE is fully compatible with the widely used Flash Attention [13] and some visual token pruning methods like FastV[8].
@@
  • We observe and summarize three main starges of MLLMs during inference, and reveal the dependency between text token status and visual redundancy, which are generalized across models and tasks.
  • Based on these empirical findings, we propose a simple yet effective method to address the visual redundancy of MLLMs, termed dynamic visual-token exit (DyVTE).
  • The extensive experiments on a set of MLLMs well validate the motivation and effectiveness of DyVTE, facilitating the future research of MLLMs.
@@ 2 Related Work
Driven by the success of large language models (LLMs) [1, 19, 2, 17, 28], the decoder-only paradigm model is also used as the main modeling approach of multimodal large language models (MLLMs) [34, 40, 3, 33, 12]. In MLLM, visual information is added mainly by using visual tokens as an extension of the input sequence. For instance, BLIP-2 [34] introduces QFormer which alleviates the modality gap while taking the most relevant content of the text as the extension of the input sequence. Qwen-VL [3] uses learnable tokens and cross-attention to collect information from image features and expand the input sequence. LLaVA [40] uses a projection layer to map visual features into the semantic space of the LLM, and take all visual tokens to expand the input sequence. Recently, in order to achieve better performance, MLLMs have invariably increased the resolution of the input image. It significantly increases the length of the input sequence. For example, LLaVA-OneVision [33] improves the resolution of the input image by dividing the high-resolution image into several regions and extracting features separately. InternVL-1.5 [12] not only introduces a stronger visual encoder, but also sets different input resolutions and scales the input image to the closest aspect ratio as the input of the visual encoder. Similarly, Qwen2-VL [53] shares the dynamic resolution with InternVL-1.5, and also proposes a carefully designed Position Embedding for various input formats. In summary, increasing resolution brings significant performance improvements, but also a significant increase in computational overhead.
To maintain the powerful capabilities of MLLMs while enhancing their efficiency, several recent works have focused on improving the computational efficiency of MLLMs. These methods, primarily in transformer-based networks [51, 15, 18], can be categorized into three main approaches: model architecture [36, 55, 56, 23, 61], dynamic reasoning [20, 21, 47, 38, 14], and token pruning [58, 29, 44, 43]. Works in model architecture [36, 55, 56] improve the efficiency of MLLMs by sparsifying the model structure. For instance, MoE-LLaVA [36] selectively activates the most relevant feedforward networks (FFNs) from a set of parallel FFNs, treating the selected FFN as the expert for the input. Similarly, RoE [55] models an MLLM as a mixture-of-experts (MoE) model, where only a subset of layers is activated, constructing a more effective and efficient expert model for each input. In the dynamic reasoning approach [21, 10], efficiency gains are achieved by terminating inference early, i.e., skipping the remaining layers. AdaInfer [21], for example, predicts the confidence of decoding tokens and terminates inference when sufficient confidence is reached. Beyond top-1 [38] analyzes the decoding token determination rules in LLMs, halting reasoning once the second-most confident token appears, thus reducing unnecessary computation. Token pruning methods [30, 54, 16, 22, 5, 52] improve reasoning efficiency by removing task-irrelevant visual tokens. FastV [8] prunes tokens based on the average attention each token receives, eliminating those deemed less important. FitPrune [59] maintains the original attention distribution during pruning, ensuring the MLLMs’ attention mechanisms remain intact. In contrast to these approaches, we propose a simple yet effective method called dynamic visual-token exit (DyVTE). By observing the pattern of MLLMs, DyVTE removes visual tokens when they no longer contribute meaningfully to multimodal reasoning, optimizing efficiency while maintaining model performance.
@@ 3.1 Method
In this paper, we propose a simple yet effective method to reduce the visual redundancy of MLLMs, termed dynamic visual-token exit (DyVTE). As discussed above, DyVTE uses lightweight hyper-networks to perceive the text token statuses of MLLMs, thereby judging the right time to remove all visual tokens to improve efficiency.
Figure 2: Illustration of the proposed dynamic visual-token exit (DyVTE). Our DyVTE aims to early exit visual tokens according to the prediction of a lightweight hyper-network based on text token status, thus achieving better efficiency for MLLMs.
As shown in Fig.2, given the text tokens $\mathbf{T}^{(k)}$ at the $k$-th layer of an MLLM, we use a simple MLP as the hyper-network to learn their statuses and decide whether to remove all visual tokens, defined by
@@
where $\mathbf{p}\in\mathcal{R}^{2}$ is the binary prediction, $\mathbf{W}_{1}\in\mathcal{R}^{2d\times h}$ and $\mathbf{W}_{2}\in\mathcal{R}^{h\times 2}$ are weight matrices, and $h$ is the dimension of the hidden states. Here, the hyper-network uses two token representations, i.e., $avg(\mathbf{T}_{1:t-1})$ and $\mathbf{T}_{t}$. The former represents the average state of most text tokens. Besides, we also use the last text token $\mathbf{T}^{(k)}_{t}$ as reference, since it often plays an important role in decoding under the uni-directional self-attention operation of MLLMs [39, 47].
When the prediction $\mathbf{p}$ refers to early visual-token exit at $l$-th layer, DyVTE will remove all visual tokens at this layer, while keeping the text tokens continues to transform in MLLMs. Given an input sample $[\mathbf{V},\mathbf{T}]$, DyVTE first gain the text tokens $\mathbf{T}^{(l)}$ and visual tokens $\mathbf{V}^{(l)}$ at $l$-th layer:
$=G_{1:l}([\mathbf{V},\mathbf{T}]),$ (2)
where $G_{1:l}(\cdot)$ represents the transformation of the first $l$ layers in the MLLM. Then all visual tokens are removed, and the final output $P^{\prime}_{l}$ with DyVTE is obtained by
@@
where $G_{l+1:L}(\cdot)$ denote the remaining layers in the MLLM.
In particular, given sufficient early-exit examples, DyVTE can learn to judge the text token statues at each layer of the MLLM. Thus, its accurate removals of visual tokens will greatly speed up the inference in the rest layers of the MLLM, while retaining similar predictions.

### v2

# Accelerating Multimodal Large Language Models via Dynamic Visual-Token Exit and Empirical Findings

@@ Abstract
In this paper, we study the visual redundancy problem of multimodal large language models (MLLMs) from the perspective of attention behaviors. Via extensive empirical experiments, we observe and conclude three main inference stages of MLLMs: (i) Early fusion between tokens is first accomplished quickly. (ii) Intra-modality modeling then comes to play. (iii) Multimodal reasoning resumes and lasts until the end of inference. In particular, we reveal that visual tokens will stop contributing to reasoning when the text tokens receive enough image information. Based on this observation, we propose an effective method to improve the efficiency of MLLMs, termed dynamic visual-token exit (DyVTE), which is orthogonal but collaborative to previous token-wise visual compression methods. To validate the efficacy of DyVTE, we apply it to a set of MLLMs, including LLaVA, VILA, EAGLE and InternVL. The experimental results not only show the effectiveness of our DyVTE in improving MLLMs’ efficiency, e.g., DyVTE reduces the computation overhead of LLaVA-1.5 by up to 45.7% without performance drop, but also reveal a general pattern across multiple MLLMs, well facilitating the in-depth analysis of MLLMs. Our code is released at https://github.com/DoubtedSteam/DyVTE.
@@ 1 Introduction
Recently, the rapid development of vision-language learning has been witnessed with the great success of large language models (LLMs) qwen ; gpt3 ; llama3 ; reid2024gemini ; qwen2.5 ; touvron2023llama ; llama2 . Numerous efforts are devoted to equip LLMs with multimodal capability, i.e., multimodal large language models (MLLMs) internvl1.5 ; internlm ; llava-onevision ; blip2 ; VILA ; llava-1.5 ; qwenvl2 . To overcome visual shortcoming pope ; textvqa in extreme high computation overhead, recent MLLMs often resort to higher-resolution images as input, which is often accompanied with a multitude of visual tokens internvl1.5 ; dong2025internlm ; llava-onevision ; luo2024feast .
@@
However, the multitude of visual tokens lead to prohibitively expensive computation. For instance, compared with LLaVA 1.5 llava-1.5 , LLaVA-NeXT llava-onevision adopts about 1728 more visual tokens, resulting in about 4 times more FLOPs. Despite the performance gains, recent worksFitPrune ; FastV ; lin2024boosting also show that these large amount of tokens are obviously redundant, leading to a great waste of computation. For instance, Ye et al. FitPrune show that randomly dropping half visual tokens barely affects performance on common VL tasks, such as MMB mmbench and SQA sqa-i . In this case, the efficient use of visual tokens has recently become a research hot-spot, and attracts an influx of attention chen2023diffrate ; dong2023heatvit ; wang2024zero ; wei2023joint . However, existing research mainly focuses on evaluating token-wise redundancy, e.g., token pruning FastV ; FitPrune and token merging li2024tokenpacker ; prumerge , lacking in-depth exploration of the intrinsic behaviors of MLLMs.
@@
In this paper, we are dedicated to understand how MLLMs use visual tokens and how they behave during multimodal reasoning. To approach this target, we conduct extensive empirical studies about the attention behaviors of MLLMs, and observe and summarize three main stages of their inference, as illustrated in Fig.1-Left. At the first stage, an MLLM quickly accomplishes the exchange of multimodal information at its shallow layers, which we term (i) early fusion. Afterwards, this information exchange decreases, and the tokens primarily engage in intra-modal interactions, termed (ii) intra-modality modeling. At the last stage, visual tokens will resume the information propagation to the text ones, and this process will continue until a certain layer, and we term it (iii) multimodal reasoning. Through extensive comparisons, we reveal that these observed behaviors of MLLMs are common across different models InternVL1 ; VILA ; llava-1.5 ; EAGLE and tasks vqav2 ; mmbench ; mm-vet .
Behind these shared behaviors of MLLMs, we observe a key finding in terms of visual redundancy, that is visual tokens will continue to propagate their semantics to the text ones at the multimodal reasoning stage, while their impact can be very limited. As shown in the right sub-figure of Fig.1, removing visual tokens at a suitable layer of MLLMs does not produce significant changes to the text self-attention distributions. This case suggests that visual tokens actually contribute little to multimodal reasoning at the last inference stage. In other words, multimodal reasoning only happens within text tokens after they receive enough visual semantics. This finding is also confirmed in some very recent works lin2024boosting , where the manual removal of visual tokens does not decline the performance of MLLMs on some tasks. But in this paper, we also found that the optimal time to remove visual tokens is distinct for different examples, tasks and even MLLMs. And the manual exploration is experimentally expensive and hard to meet the global optimum. Therefore, when and how to automatically remove visual tokens is still a thorny challenge.
Figure 1: Left: Illustration of three main stages observed in MLLMs. During inference, an MLLM will go through three main stages, i.e., early fusion between all tokens, intra modality modeling of the same-modal tokens, and multimodal reasoning between all tokens again. Right: The impact of visual tokens on the text self attention at the multimodal reasoning stage. Removing visual tokens at an appropriate time barely changes the text attention patterns, indicating that visual tokens contribute little to multi-modal reasoning.
@@
Motivated by these observations, we propose a novel and effective method to improve MLLMs’ efficiency, termed dynamic visual-token exit (DyVTE). In particular, we apply lightweight hyper-networks to perceive text token status and then dynamically decide the exit of all visual tokens, thereby speeding up inference. As discussed above, DyVTE mainly focuses on the overall contribution of all visual tokens during multimodal inference, which is orthogonal but collaborative to token-wise solutions like token pruning FastV ; kong2022spvit ; FitPrune . In our experiments, we show that these two paradigms can be easily combined to boost the efficiency of MLLMs. In addition, DyVTE also differs from previous dynamic early exiting methods fei2022deecap ; ghodrati2021frameexit ; tang2023you , which often refer to an inference break i.e., skipping rest layers for direct predictions. In contrast, text tokens continue to transform in DyVTE, which better meets our empirical findings. Moreover, we also introduce an efficient training regime for DyVTE independent to the global gradient computation of MLLMs, achieving effective and dynamic visual-token exit with very cheap training expenditure.
@@
To validate DyVTE, we apply it to a set of advanced MLLMs with varying scales, including LLaVA-1.5 llava-1.5 , VILA VILA , EAGLE EAGLE and InternVL InternVL1 , and conduct experiments on a bunch of widely-used VL and MLLM benchmarks mme ; vqav2 ; gqa ; pope ; textvqa ; mm-vet . The experimental results show that our DyVTE can greatly reduce the computation overhead of MLLMs, while retaining their competitive performance on various benchmarks. For instance, our DyVTE reduces the computation overhead of LLaVA-1.5 by up to 45.7% without performance drop. When combined with token pruning methods like FastV FastV , DyVTE can help LLaVA-1.5 reduce up to 51.5% computation with only 0.7% performance drop on average.
@@
  • We study the problem of visual redundancy from the perspective of MLLMs’ behaviors, and reveal the dependency between text and visual tokens.
  • Based on the empirical findings, we propose a novel and effective approach to reduce visual redundancy of MLLMs, termed dynamic visual-token exit (DyVTE), which can dynamically evaluate and schedule the contributions of visual tokens to multimodal reasoning.
  • The extensive experiments on a set of MLLMs well validate the motivation and effectiveness of DyVTE, also providing insights into the principle of MLLMs.
@@ 2 Related Work
Based on the successful large language models (LLMs) gpt4 ; llama3 ; qwen ; internlm ; mixtral , numerous multimodal large language models (MLLMs) have been recently proposed blip2 ; llava-1.5 ; qwenvl ; llava-onevision ; internvl1.5 and achieved remarkable progresses on various vision-language tasks vqav2 ; mme ; gqa . In terms of methodology, most MLLMs often project the extracted visual features onto the semantic space of LLM for both multimodal fusion and reasoning llava-1.5 ; qwenvl2 ; EAGLE ; InternVL1 . For instance, LLaVA llava-1.5 uses a projection layer to transform visual features into input tokens for LLaMA llama2 . BLIP-2 blip2 introduces QFormer to aggregate query-related visual features as the input visual tokens. Similarly, QWen-VL also adopts learnable tokens as queries to obtain a limited number of visual tokens qwenvl . To improve the ability in fine-grained tasks textvqa ; infovqa ; docvqa , some recent MLLMs resort to increasing the resolution of input images llava-onevision ; qwenvl2 ; internvl2.5 for better visual understanding. For instance, LLaVA-NeXT llava-next concatenates the tokens of each crop of the image as the input of the LLM. InternVL2.5 internvl2.5 and Qwen2-VL qwen2.5vl introduce dynamic resolution designs to match different images and preserve their visual details. Despite success, these high-resolution settings also inevitably increase the number of visual tokens, leading to excessive computation overhead FastV ; FitPrune .
Moreover, recent studies also show that the excessive use of visual tokens brings in obvious redundancy in both information and computation moe-llava ; roe ; das . In this case, the efficiency study of MLLMs also becomes a research hotspot, of which focus ranges from structure design moe-llava ; roe ; das , token pruning kong2022spvit ; fayyaz2022adaptive ; ToMe and token merging prumerge ; ToMe . The research of structure design mainly aim to build a new and lightweight MLLMs moe-llava ; tinyllava , and our work is closer to the token efficiency researches FitPrune ; FastV . For instance, FastV FastV applies the averaged attention scores to evaluate the importance of each visual token and then drop the less important ones to reduce computation. FitPrune FitPrune resort to a statistic principle to quickly produces a token pruning regime for MLLMs based on the metrics of visual and cross attentions. PruMerge prumerge speeds up the inference of MLLMs by merging visual tokens that have similar semantics. G-Prune jiang2025kind proposes an graph-based algorithm to select the most representative visual tokens to avoid redundant semantics and computations. Compared with these token-wise approaches, our DyVTE focuses more on the overall contributions of visual tokens in MLLMs, providing an alternative but orthogonal way for efficient MLLMs. In terms of dynamic inference, our works are also related to the dynamic early-exit studies for LLMs AdaInfer ; ee-llm . However, these methods focus on the skipping of redundant layers for early prediction akin to previous dynamic methods tang2023you ; ghodrati2021frameexit . Recently, LLaVA-Mini LLaVA-mini introduces a pre-fusion module through full pre-training and SFT, reducing visual tokens to just one. And our DyVTE is to conduct the early removal of visual tokens while keeping the original inference pattern. Moreover, this paper also involves quantitative analyses about the attention behaviors of MLLMs, providing insights into the mechanisms for their multimodal reasoning.
3 Inference Pattern Analysis of MLLMs
We first quantitatively investigate the attention behaviors of MLLMs, and then examine the impacts of visual tokens during multimodal inference.
Figure 2: The averaged attention scores of four MLLMs in terms of cross, visual and text attentions. The background colors denote the three summarized stages of MLLMs, i.e., early fusion (green), intra-modality modeling (purple) and multimodal reasoning (blue).
Figure 3: The relationship between manual visual-token exit and performance on LLaVA-1.5 7B and 13B. We show the results of removing all visual tokens at 0-th, 888-th, 161616-th, 242424-th and 323232-th layers. “Baseline” represents the original performance of the MLLM. After a certain layer, the removal of all visual tokens barely impedes performance, but the layers for removal are different for different tasks.
Figure 4: The entropy of text self-attention distribution with different layer to remove all visual tokens. We visualize the entropy of text self-attention on LLaVA-1.5 7B on GQA benchmark with different with different layer to remove all visual tokens, which shows that the removal of visual tokens after a certain layer barely affect the text self-attention.
Attention behaviors of MLLMs. We first visualize the attention patterns of four representative MLLMs llava-1.5 ; EAGLE ; InternVL1 of two scales in Fig.2. These attention data includes visual self-attention, visual-text cross-attention and text self-attention [^1] The visualization details are provided in our appendix A.1.. Here, the values in these distributions are the column-wise summarized attention scores on the LLaVA-split llava-1.5 . From these plots, we can first observe that four MLLMs share similar patterns for three types of attentions. At the shallow layers, all MLLMs show high cross-attention values between visual and text tokens, suggesting the intensive interactions between two modalities. This corresponds to the early fusion stage we termed above, and it denotes the quick exchange of multimodal information. But early fusion declines quickly, and we can observe that the self-attention between the tokens of the same modality becomes the main activity at the second stage, i.e., the modeling mainly involved in each modality. In this case, we can assume that MLLM starts the intra-modality modeling stage. At the last stage, the visual tokens will resume information propagation to the text tokens, and this process will continue to oscillate until the end of inference. Here, we term this process as the multimodal reasoning stage. Notably, we can find these patterns exist in MLLMs of different families and scales.
[…]
Discussion. Based on the above observations and analyses, we can categorize the inference of MLLMs into three main stages. (i) Early Fusion. In the initial stage, the MLLM quickly accomplish the exchange of multimodal information from visual to text tokens. (ii) Intra-modality Modeling. Next, the self-attention intra the modality becomes the main activity, enhancing vision and language understanding. (iii) Multimodal Reasoning. Finally, the visual tokens will resume the information propagation to the text tokens, though this transfer demonstrates limited impact on final response generation. In addition to these three stages, we can also conclude one common behavior of visual tokens in MLLMs, i.e., the contribution of visual tokens to multimodal prediction diminish after some layers. Moreover, these results also show that the exist times for different types of VL examples are different, well motivating our exploration of dynamic visual-token exit.
@@ 4.1 Method
In this paper, we propose a novel and effective method to reduce the visual redundancy of MLLMs, termed dynamic visual-token exit (DyVTE). DyVTE uses lightweight hyper-networks to perceive the text token status of MLLMs and then adaptively judge the right time to remove all visual tokens.
Concretely, given the text tokens $\mathbf{T}^{(k)}$ at the $k$-th layer of an MLLM, we use a simple MLP as the hyper-network to learn their status and decide whether to remove all visual tokens, defined by
@@
where $\mathbf{p}\in\mathcal{R}^{2}$ is the binary prediction, $\mathbf{W}_{1}\in\mathcal{R}^{2d\times h}$ and $\mathbf{W}_{2}\in\mathcal{R}^{h\times 2}$ are weight matrices, and $h$ is the dimension of the hidden states.
In particular, the hyper-network of Eq.1 uses two text token representations, i.e., $avg(\mathbf{T}_{1:t-1})$ and $\mathbf{T}_{t}$. The former is the averaged text tokens representing their overall state. The latter $\textbf{T}_{t}$ is the last token, which often plays an important role for decoding under the uni-directional self-attention setting of MLLMs lioubashevski2024looking ; stolfo2023mechanistic . With these two types of text representations, DyVTE can well perceive the state of MLLMs and automatically decide whether additional visual information is still needed for reasoning.
When the prediction $\mathbf{p}$ is exit, DyVTE will remove all visual tokens after this layer, while the text ones are kept in the rest layers of MLLMs. This process can be denoted by
@@
where $G_{l+1:L}(\cdot)$ denote the remaining layers in the MLLM. Compared with previous dynamic exit methods that skip layers for early prediction AdaInfer ; ee-llm , the principle of our DyVTE is to make the removal of visual tokens based on the text token status. Thus, its implementation requires no structure tweaks, e.g., adding new prediction layers, and also no the update of MLLMs.